Bus fares from Amherst to Mount Mohawk are $26 for adults and $18 for students. How many students are on the bus if a total of $1668 was collected from the 70 passengers?
Answer by checkley71(8403)

26X+18(70-X)=1668
26X+1260-18X=1668
8X=1668-1260
8X=408
X=408/8
X=51 NUMBER OF ADULT TICKETS SOLD.
70-51=19 STUDENT TICKETS SOLD.
PROOF
26*51+18*19=1668
1326+342=1668
1668=1668